OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)] −13 −25 25 17

OpenStudy (anonymous):

I nhink -13

OpenStudy (anonymous):

f(g(-2))=f(-13)=25

hero (hero):

f(x) = -2x - 1 g(x) = 6x - 1 f(g(x)) = -2(6x - 1) - 1 f(g(x)) = -12x +2 - 1 = -12x + 1 f(g(-2)) = -12(-2) + 1 = 24 + 1 = 25

OpenStudy (lalaly):

\[f(g(x))=-2(6x-1)-1=-12x+1\]\[f(g(-2)=-12(-2)+1=24+1=25\]

OpenStudy (anonymous):

:)

OpenStudy (lalaly):

lol Thanks ghassan:)

OpenStudy (anonymous):

yw,i'm the one givin medals not the asker:)